OpenStudy (anonymous):

Find the area of a parallelogram with the given vertices. P(1, 3), Q(3, 3), R(7, 8), S(9, 8) A. 10 units2 B. 5 units2 C. 20 units2 D. none of these

OpenStudy (anonymous):

@Nurali @wolf1728

OpenStudy (nurali):

The length of a horizontal side is PQ or RS, which has a length of 3-1 = 2 The height of the parallelogram is 8-3 = 5 (y direction difference) Then the area is A = BH (base times height) A = 2*5 = 10

OpenStudy (anonymous):

Thanks again